OpenStudy (anonymous):

A 20 pound bag of sugar costs (y + 3) dollars, while a 10 pound package costs (y - 3) dollars. If the price per pound of sugar is the same for both packages, the equation below can be used to solve for y. what is the value, in dollars, of y? (a)y = 3 (b) y = 9 (c) y = 30 (d) y = 90

OpenStudy (4n1m0s1ty):

Price per pound must be the same, therefore: (y+3)/20 = (y-3)/10 Get the y's on one side: Multiply by 20 y+3 = 2(y-3) Distribute right side y+3 = 2y - 6 add 6 y+9 = 2y subtract y y = 9

OpenStudy (anonymous):

thank youuuu :)